How much water should be added to 1 gallon of pure antifreeze to obtain a solution that is 60% antifreeze?
Answer by Alan3354(21605)

How much water should be added to 1 gallon of pure antifreeze to obtain a solution that is 60% antifreeze?
---------------
The original 1 gallon of a/f will be 60% of the solution, s.
1 gallon = 60% = 0.6s
s = 1/0.6 = 5/3 gallons
Add 2/3 gallons.